# Conservative Perception Models for Probabilistic Verification

### v2

@@ 6. Theoretical Guarantee of Conservative Verification
Let bin $s_{i}\in S$ denote a discrete perception bin with corresponding state space region $\mathcal{X}_{i}\subseteq\mathcal{X}$, as described in Section 5.1. Now let $P_{s_{i}}(h)$ denote the range of the detection probabilities of perception $h$ over the states $x\in s_{i}$:
(7) $P_{s_{i}}(h)=\left[\min_{x\in s_{i}}P_{x}(h),\max_{x\in s_{i}}P_{x}(h)\right]$
@@
Consider the same inputs as in Problem 1. Let $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ denote the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval of the perception component detection probability in bin $s_{i}$ (as described in Section 5.2). Assume [^5] In practice, this assumption means that the true detection probabilities can be instantiated by some parameter values of Step 3’s surrogate model, e.g., the logistic regression. In our case studies, this assumption has been largely upheld. Its theoretical investigation is left for future work. the probability enlargement factor $\delta_{s_{i}}$ as computed in Equation 4 upper-bounds the size of $P_{s_{i}}(h)$:
(8) $\delta_{s_{i}}\geq\left|\max_{x\in s_{i}}P_{x}(h)-\min_{x\in s_{i}}P_{x}(h)\right|$
@@
$\underline{\hat{p}}_{s_{i}}-\delta^{p}_{s_{i}}\leq\min_{x\in s_{i}}P_{x}(h)$
$\bar{\hat{p}}_{s_{i}}+\delta^{p}_{s_{i}}\geq\max_{x\in s_{i}}P_{x}(h)$
@@
(9) $P_{\mathsf{M}_{\text{sys}}}(\psi)\in[P_{\mathsf{M}_{\text{abs}}}^{min}(\psi),P_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]$
7. Case Studies
Our experimental evaluation has two goals: explore how the bin size hyperparameter and how the probability enlargement step affect the conservatism of our approach on a fully synthetic case study. We use an automatic emergency braking system (AEBS) on an autonomous car to evaluate our approach. Additionally, we apply our approach to an autonomous car example in the CARLA simulator (dosovitskiy_carla_2017, ). For IMDP verification, we use PRISM model checker 4.8.1 with native IMDP support.
7.1. System: Emergency Braking with Object Detection
@@
(10) $d[t]=d[t-\tau]-\tau\times v[t-\tau],\quad v[t]=v[t-\tau]-\tau\times b[t-\tau],$
@@
We use an Advanced Emergency Braking System (AEBS) (lee_aebs_2011, ) which uses two metrics to determine the BP: time to collide ($TTC$) and warning index ($WI$). The $TTC$ is the amount of time until a collision if the current velocity is maintained. The $WI$ represents how safe the car would be in the hands of a human driver (positive is safe, negative is unsafe).
(11) $TTC=\frac{d}{v},\quad WI=\frac{d-d_{br}}{vT_{h}},\quad d_{br}=vT_{s}+\frac{uv^{2}}{2a_{max}},$
where $d_{br}$ is the braking-critical distance, $T_{s}$ is the system response delay (negligible), $T_{h}$ is the average driver reaction time (set to 2s), $u$ is the friction scaling coefficient (taken as 1), and $b_{max}$ is the maximum deceleration of the car.
@@
$WI>C_{1}\land TTC>C_{2}\implies$ $b=0$
$(WI\leq C_{1}\land TTC>C_{2})\lor(WI>C_{1}\land TTC\leq C_{2})\implies$ $b=B_{1}$
$WI\leq C_{1}\land TTC\leq C_{2}\implies$ $b=B_{2}$
To detect the obstacle, the car uses a classification variant of the deep neural network YoloNetv11 (yolo11_ultralytics, ), as it can run at high frequencies. We say a low-level detection of the obstacle occurs when Yolo detects the obstacle.
@@
(12) $\psi_{nocol}:=\square\;\left(d>L\right),$
@@
7.2. Comparative Abstraction Methods
We compare our approach to a few baseline/ablation methods of creating $\textsf{M}_{\text{per}}$ summarized in Table 1: a baseline approach (badithela_evaluation_2023, ) that doesn’t use confidence intervals, our approach with no probability enlargement (this is our approach, but skipping Section 5.3), and an approach that uses confidence intervals from the logistic regression we employ in Section 5.3 to compute the bin probability intervals, and the approach that uses the ground truth perception detections probabilities (for the synthetic case study only). We expect our approach to be more conservative than all of these due to its guarantees.
@@ 7.2. Comparative Abstraction Methods
$\mathsf{M}_{noCI}$ Baseline using $\bar{p}_{s_{i}}$ from Equation 2 as detection probabilities (badithela_evaluation_2023, )
@@
Table 1. Perception abstraction approaches compared in our experiments.
7.3. Synthetic perception case study
@@
Figure 3. Detection probabilities for the synthetic perception model.
@@ 7.3.1. Bin Width Effect:
Overall, this experiment highlights that our approach produces conservative models regardless of bin size — while the existing approaches do not.
@@
Figure 4. Results of the synthetic perception case study.
7.3.2. Probability Enlargement Factor:
@@
(14) $[\underline{\hat{p}}_{s_{i}}-w_{PE}\delta^{p}_{s_{i}},\bar{\hat{p}}_{s_{i}}+w_{PE}\delta^{p}_{s_{i}}]$
Using the same experimental conditions as before, we tested our approach using $w_{PE}\in\{0,0.3,0.7,1\}$, where $w_{PE}=1$ corresponds to our full approach $\mathsf{M}_{ours}$, and $w_{PE}=0$ corresponds to the no probability enlargement $\mathsf{M}_{oursNPE}$. The results of this experiment are shown in Figure 4(b). First, note that the safety intervals are monotonically increasing as the factor increases. That is because the larger probability enlargement factor increases the detection probability intervals in $\textsf{M}_{\text{per}}$. For $w_{PE}=0.7$, the approach remains empirically conservative. We conclude that bridging the gap between the probability enlargement needed for theoretical and empirical conservatism is an interesting avenue for future research.
7.4. CARLA case study
Here, we evaluate the feasibility of our approach for a realistic object detection LEC in a high-fidelity simulation. First, we set up the AEBS controller in the autonomous driving simulator CARLA (dosovitskiy_carla_2017, ), using a red Toyota Prius as the obstacle and a classification variant of Yolov11 as the object detector. Then we performed two rounds of data collection in CARLA: (i) constant-$v$ runs to gather image-distance data to build our perception abstractions, (ii) AEBS-controlled runs to approximate the true safety probability of the AEBS controller.
For the constant-$v$ runs, we ran $500$ trials with a speed of $6$m/s and initial distance around $210$m, which generated $181000$ perception datapoints. Figure 5(a) shows 5m-wide binned estimates of the YoloNet detection probabilities from the perception data. For the AEBS-controlled runs, we ran $500$ trials with an initial distance $50$m and an initial speed $20$m/s. This resulted in $330$ crashes, which yields a $[0.62,0.70]$ 95%-CI for the true collision chance.
@@ 7.4. CARLA case study
Figure 5. Results of the CARLA case study.
We used the perception data to construct models using our approach $\mathsf{M}_{ours}$, our approach with no probability enlargement $\mathsf{M}_{oursNPE}$, the approach from (badithela_evaluation_2023, ) with no confidence intervals $\mathsf{M}_{noCI}$, and the approach using the logistic regression confidence intervals $\mathsf{M}_{logRegCI}$. Note that we cannot construct $\mathsf{M}_{GTPer}$ since we don’t have access to the ground truth detection probabilities.
We first vary the bin width parameter used to form the models to determine how it affects the estimated safety probability intervals. The results are shown in Figure 5(b). Note that for every bin width the safety probability intervals from our approach contain the confidence interval for the true underlying safety probability. Also, the probabilities from $\mathsf{M}_{oursNPE}$ are only empirically valid for small bins, as we witnessed in Section 7.3. Additionally, the probabilities from $\mathsf{M}_{noCI}$ never contain the true safety probability. Finally, the safety probabilities from $\mathsf{M}_{logRegCI}$ are not valid for small bin widths. This highlights the drawbacks of solely relying on logistic regression to estimate detection probabilities in practice.
8. Conclusion and Future Work
@@
Future work can focus on reducing conservatism by better estimating intra-bin changes in detection probabilities of the perception component. For example, more sophisticated surrogate models for probability
@@ A.1. Proof of Lemma 3
Consider the same inputs as in Problem 1. Let $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ denote the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval of the perception component detection probability in bin $s_{i}$ (as described in Section 5.2). Assume the probability enlargement factor $\delta_{s_{i}}$ as computed in Equation 4 upper-bounds the size of $P_{s_{i}}(h)$:
(15) $\delta_{s_{i}}\geq\left|\max_{x\in s_{i}}P_{x}(h)-\min_{x\in s_{i}}P_{x}(h)\right|$
$\underline{\hat{p}}_{s_{i}}-\delta^{p}_{s_{i}}\leq\min_{x\in s_{i}}P_{x}(h)$
$\bar{\hat{p}}_{s_{i}}+\delta^{p}_{s_{i}}\geq\max_{x\in s_{i}}P_{x}(h)$
@@
By the use of the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval when computing $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$, we know that average detection probability of $\textsf{M}_{\text{per}}$ over bin $s_{i}$ is contained in $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ with probability $1-\frac{\alpha_{MC}}{|S|}$. Next, by the assumption that $\delta_{s_{i}}$ upper bounds the size of $P_{\textsf{M}_{\text{per}}}(s_{i})$, it follows that
$\underline{\hat{p}}_{s_{i}}-\delta^{p}_{s_{i}}\leq\min_{x\in s_{i}}P_{\textsf{M}_{\text{per}}}(x)$
$\bar{\hat{p}}_{s_{i}}+\delta^{p}_{s_{i}}\geq\max_{x\in s_{i}}P_{\textsf{M}_{\text{per}}}(x)$
A.2. Proof of Theorem 4
@@
(16) $P_{\mathsf{M}_{\text{sys}}}(\psi)\in[P_{\mathsf{M}_{\text{abs}}}^{min}(\psi),P_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]$
@@
Let $\mathsf{M}_{\text{sys}}$ be in state $x\in\mathcal{X}$. Its perception component $h$ either outputs a detection ($h(g(x,v)=1$) with probability $P_{x}(h)$ or a non-detection ($h(g(x,v)=0$) with probability $1-P_{x}(h)$ (recall that $v$ is an unknown, random noise variable). In the event of a detection, the next state of $\mathsf{M}_{\text{sys}}$ will be $x_{1}^{\prime}=f(x,c(y,1))$ and in the event of a non-detection, the next state of $\mathsf{M}_{\text{sys}}$ will be $x_{0}^{\prime}=f(x,c(y,0))$. This induces the distribution of the next state $x^{\prime}$ of $\mathsf{M}_{\text{sys}}$.
Now consider our abstraction $\mathsf{M}_{\text{abs}}$ stating in its state $s$ that corresponds to $x$ (i.e. $x\in s$). Let’s assume that the perception bin detection probability interval of $\textsf{M}_{\text{per}}$ for state $x$ contains $P_{x}(h)$. So the probability interval that $\textsf{M}_{\text{per}}$ sends $a_{per}=1$ to $\mathsf{M}_{\text{cpl}}$ contains the true detection probability $P_{x}(h)$ and the probability interval that $\textsf{M}_{\text{per}}$ sends $a_{per}=0$ contains the true mis-detection probability $1-P_{x}(h)$. By the assumption that $\mathsf{M}_{\text{cpl}}$ is conservative (per Problem 1), if $\textsf{M}_{\text{per}}$ sends a detection ($a_{per}=1$) then the set of next states $S_{1}^{\prime}=\{s^{\prime}\in S|\exists a_{reach}\in\alpha_{reach}\text{ suchthat }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{1,a_{reach}\},s^{\prime})=1\}$ that $\mathsf{M}_{\text{cpl}}$ can transition to will contain $x_{1}^{\prime}$ ($\exists s_{1}^{\prime}\in S_{1}^{\prime}$ such that $x_{1}^{\prime}\in s_{1}^{\prime}$) and if $\textsf{M}_{\text{per}}$ sends a no detection ($a_{per}=0$) then the set of next states $S_{0}^{\prime}=\{s^{\prime}\in S|\exists a_{reach}\in\alpha_{reach}\text{ suchthat }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{0,a_{reach}\},s^{\prime})=1\}$ that $\mathsf{M}_{\text{cpl}}$ can transition to will contain $x_{0}^{\prime}$ ($\exists s_{0}^{\prime}\in S_{0}^{\prime}$ such that $x_{0}^{\prime}\in s_{0}^{\prime}$). So starting from state $x$, the distribution of the next state defined by $\mathsf{M}_{\text{cpl}}||\textsf{M}_{\text{per}}$ contains the distribution of the next state for $\mathsf{M}_{\text{sys}}$ if the detection probability intervals for every bin of $\textsf{M}_{\text{per}}$.
@@
Finally, noting that $\textsf{M}_{\text{per}}$ has $|S|$ bins, the probability that every perception bin’s detection probability interval contains the true range of detection probabilities is
$\sum_{b\in S}1-\frac{\alpha_{MC}}{|S|}\leq 1-\alpha_{MC}$

### v3

@@ VI Conservative Verification Guarantee
Let bin $s_{i}\in S$ denote a discrete perception bin with corresponding state space region $\mathcal{X}_{i}\subseteq\mathcal{X}$, as described in Section V-A. Now let $P_{s_{i}}(h)$ denote the range of the detection probabilities of perception $h$ over the states $x\in s_{i}$:
$P_{s_{i}}(h)=\left[\min_{x\in s_{i}}P_{x}(h),\max_{x\in s_{i}}P_{x}(h)\right]$ (17)
@@
Consider the same inputs as in Problem 1. Let $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ denote the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval of the perception component detection probability in bin $s_{i}$ (as described in Section V-B). Assume [^5] In practice, this assumption means that the true detection probabilities can be instantiated by some parameter values of Step 3’s surrogate model, e.g., the logistic regression. In our case studies, this assumption has been largely upheld. Its theoretical investigation is left for future work. the probability enlargement factor $\delta_{s_{i}}$ as computed in Equation 14 upper-bounds the size of $P_{s_{i}}(h)$:
$\delta_{s_{i}}\geq\left|\max_{x\in s_{i}}P_{x}(h)-\min_{x\in s_{i}}P_{x}(h)\right|$ (18)
@@
$\underline{\hat{p}}_{s_{i}}-\delta_{s_{i}}\leq\min_{x\in s_{i}}P_{x}(h)$
$\bar{\hat{p}}_{s_{i}}+\delta_{s_{i}}\geq\max_{x\in s_{i}}P_{x}(h)$
holds with probability $1-\frac{\alpha_{MC}}{|S|}$.
@@
$P\big{(}P_{\mathsf{M}_{\text{sys}}}(\psi)\in[Pr_{\mathsf{M}_{\text{abs}}}^{min}(\psi),Pr_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]\big{)}\geq 1-\alpha_{MC}$ (19)
VII Case Studies
Our experimental evaluation has two goals: explore how the bin size hyperparameter and how the probability enlargement step affect the conservatism of our approach on a fully synthetic case study. We use an automatic emergency braking system (AEBS) on an autonomous car to evaluate our approach. Additionally, we apply our approach to an autonomous car example in the CARLA simulator [16]. For IMDP verification, we use PRISM model checker 4.8.1 with native IMDP support.
VII-A System: Emergency Braking with Object Detection
@@
$d[t]=d[t-\tau]-\tau\times v[t-\tau],\quad v[t]=v[t-\tau]-\tau\times b[t-\tau],$ (20)
@@
We use an Advanced Emergency Braking System (AEBS) [39] which uses two metrics to determine the BP: time to collide ($TTC$) and warning index ($WI$). The $TTC$ is the amount of time until a collision if the current velocity is maintained. The $WI$ represents how safe the car would be in the hands of a human driver (positive is safe, negative is unsafe).
$TTC=\frac{d}{v},\quad WI=\frac{d-d_{br}}{vT_{h}},\quad d_{br}=vT_{s}+\frac{uv^{2}}{2a_{max}},$ (21)
where $d_{br}$ is the braking-critical distance, $T_{s}$ is the system response delay (negligible), $T_{h}$ is the average driver reaction time (set to 2s), $u$ is the friction scaling coefficient (taken as 1), and $a_{max}$ is the maximum deceleration of the car.
@@
$\begin{split}&WI>C_{1}\land TTC>C_{2}\implies b=0\\ &(WI\leq C_{1}\land TTC>C_{2})\lor(WI>C_{1}\land TTC\leq C_{2})\\ &\quad\quad\implies b=B_{1}\\ &WI\leq C_{1}\land TTC\leq C_{2}\implies b=B_{2}\end{split}$ (22)
To detect the obstacle, the car uses a classification variant of the deep neural network YoloNetv11 [28], as it can run at high frequencies. We say a low-level detection of the obstacle occurs when Yolo detects the obstacle.
@@
$\psi_{nocol}:=\square\;\left(d>L\right),$ (23)
@@
VII-B Comparative Abstraction Methods
We compare our approach to a few baseline/ablation methods of creating $\textsf{M}_{\text{per}}$ summarized in Table I: a baseline approach [6] that doesn’t use confidence intervals, our approach with no probability enlargement (this is our approach, but skipping Section V-C), and an approach that uses confidence intervals from the logistic regression we employ in Section V-C to compute the bin probability intervals (this is our approach, but skipping Section V-B), and the approach that uses the ground truth perception detections probabilities (for the synthetic case study only). We expect our approach to be more conservative than all of these due to its guarantees.
@@ VII-B Comparative Abstraction Methods
$\mathsf{M}_{noCI}$ Baseline using $\hat{p}_{s}$ from Equation 12 as detection probabilities [6]
@@
TABLE I: Perception abstraction approaches compared in our experiments.
VII-C Synthetic perception case study
@@
Figure 3: Detection probabilities for the synthetic perception model.
@@ VII-C1 Bin Width Effect:
Overall, this experiment highlights that our approach produces conservative models regardless of bin size — while the existing approaches do not.  [^7] Because our method is conservative for any bin size, the user can test a range of bin sizes and use whichever one gives the tightest safety probability interval.
@@
Figure 4: Results of the synthetic perception case study.
VII-C2 Probability Enlargement Factor:
@@
$[\underline{\hat{p}}_{s_{i}}-w_{PE}\delta^{p}_{s_{i}},\bar{\hat{p}}_{s_{i}}+w_{PE}\delta^{p}_{s_{i}}]$ (25)
Using the same experimental conditions as before, we tested our approach using $w_{PE}\in\{0,0.3,0.7,1\}$, where $w_{PE}=1$ corresponds to our full approach $\mathsf{M}_{ours}$, and $w_{PE}=0$ corresponds to the no probability enlargement $\mathsf{M}_{oursNPE}$. The results of this experiment are shown in Figure 4(b). First, note that the safety intervals are monotonically increasing as the factor increases. That is because the larger probability enlargement factor increases the detection probability intervals in $\textsf{M}_{\text{per}}$. For $w_{PE}\in\{0.3,0.7\}$, the approach remains empirically conservative. We conclude that bridging the gap between the probability enlargement needed for theoretical and empirical conservatism is an interesting avenue for future research.
VII-D CARLA case study
[…]
@@ VII-D CARLA case study
Figure 5: Results of the CARLA case study.
We used the perception data to construct models using our approach $\mathsf{M}_{ours}$, our approach with no probability enlargement $\mathsf{M}_{oursNPE}$, the approach from [6] with no confidence intervals $\mathsf{M}_{noCI}$, and the approach using the logistic regression confidence intervals $\mathsf{M}_{logRegCI}$. Note that we cannot construct $\mathsf{M}_{GTPer}$ since we don’t have access to the ground truth detection probabilities.
We first vary the bin width parameter used to form the models to determine how it affects the estimated safety probability intervals. The results are shown in Figure 5(b). Note that for every bin width the safety probability intervals from our approach contain the confidence interval for the true underlying safety probability. Also, the probabilities from $\mathsf{M}_{oursNPE}$ are only empirically valid for small bins, as we witnessed in Section VII-C. Additionally, the probabilities from $\mathsf{M}_{noCI}$ never contain the true safety probability. Finally, the safety probabilities from $\mathsf{M}_{logRegCI}$ are not valid for small bin widths. This highlights the drawbacks of solely relying on logistic regression to estimate detection probabilities in practice.
VIII Conclusion and Future Work
@@
Future work can focus on reducing conservatism by better estimating intra-bin changes in detection probabilities of the perception component. For example, more sophisticated surrogate models for probability. Additionally, we will investigate more expressive $\textsf{M}_{\text{per}}$ models, such as conditioning detections on previous $\textsf{M}_{\text{per}}$ outputs, and explore abstractions for multi-class LECs. Finally, we will extend our approach to specifications beyond LTL safety.
@@ A-A Proof of Lemma 1
Consider the same inputs as in Problem 1. Let $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ denote the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval of the perception component detection probability in bin $s_{i}$ (as described in Section V-B). Assume the probability enlargement factor $\delta_{s_{i}}$ as computed in Equation 14 upper-bounds the size of $P_{s_{i}}(h)$:
$\delta_{s_{i}}\geq\left|\max_{x\in s_{i}}P_{x}(h)-\min_{x\in s_{i}}P_{x}(h)\right|$ (26)
Then, with probability $1-\frac{\alpha_{MC}}{|S|}$, the interval from Equation 15 contains the true detection probability interval $P_{s_{i}}(h)$. Precisely:
$\underline{\hat{p}}_{s_{i}}-\delta_{s_{i}}\leq\min_{x\in s_{i}}P_{x}(h)$
$\bar{\hat{p}}_{s_{i}}+\delta_{s_{i}}\geq\max_{x\in s_{i}}P_{x}(h)$
@@
By the use of the $1-\frac{\alpha_{MC}}{|S|}$ confidence interval when computing $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$, we know that average detection probability of $\textsf{M}_{\text{per}}$ over bin $s_{i}$ is contained in $[\underline{\hat{p}}_{s_{i}},\bar{\hat{p}}_{s_{i}}]$ with probability $1-\frac{\alpha_{MC}}{|S|}$. Next, by the assumption that $\delta_{s_{i}}$ upper bounds the size of $P_{s_{i}}(h)$, it follows that
$\underline{\hat{p}}_{s_{i}}-\delta_{s_{i}}\leq\min_{x\in s_{i}}P_{x}(h)$
$\bar{\hat{p}}_{s_{i}}+\delta_{s_{i}}\geq\max_{x\in s_{i}}P_{x}(h)$
A-B Proof of Theorem 1
@@
$P_{\mathsf{M}_{\text{sys}}}(\psi)\in[Pr_{\mathsf{M}_{\text{abs}}}^{min}(\psi),Pr_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]$ (27)
@@
Let $\mathsf{M}_{\text{sys}}$ be in state $x\in\mathcal{X}$. Its perception component $h$ either outputs a detection ($h(g(x,v)=1$) with probability $P_{x}(h)$ or a non-detection ($h(g(x,v)=0$) with probability $1-P_{x}(h)$ (recall that $v$ is an unknown, random noise variable). In the event of a detection, the next state of $\mathsf{M}_{\text{sys}}$ will be $x_{1}^{\prime}=f(x,c(1))$ and in the event of a non-detection, the next state of $\mathsf{M}_{\text{sys}}$ will be $x_{0}^{\prime}=f(x,c(0))$. This induces the distribution of the next state $x^{\prime}$ of $\mathsf{M}_{\text{sys}}$.
Now consider our abstraction $\mathsf{M}_{\text{abs}}$ stating in its state $s$ that corresponds to $x$ (i.e. $x\in s$). Let’s assume that the perception bin detection probability interval of $\textsf{M}_{\text{per}}$ for state $x$ contains $P_{x}(h)$. So the probability interval that $\textsf{M}_{\text{per}}$ sends $a_{per}=1$ to $\mathsf{M}_{\text{cpl}}$ contains the true detection probability $P_{x}(h)$ and the probability interval that $\textsf{M}_{\text{per}}$ sends $a_{per}=0$ contains the true mis-detection probability $1-P_{x}(h)$. By the assumption that $\mathsf{M}_{\text{cpl}}$ is conservative (per Problem 1), if $\textsf{M}_{\text{per}}$ sends a detection ($a_{per}=1$) then the set of next states $S_{1}^{\prime}=\{s^{\prime}\in S|\exists a_{reach}\in\alpha_{reach}\text{ such that }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{1,a_{reach}\},s^{\prime})=1\}$ that $\mathsf{M}_{\text{cpl}}$ can transition to will contain $x_{1}^{\prime}$ ($\exists s_{1}^{\prime}\in S_{1}^{\prime}$ such that $x_{1}^{\prime}\in s_{1}^{\prime}$) and if $\textsf{M}_{\text{per}}$ sends a no detection ($a_{per}=0$) then the set of next states $S_{0}^{\prime}=\{s^{\prime}\in S|\exists a_{reach}\in\alpha_{reach}\text{ such that }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{0,a_{reach}\},s^{\prime})=1\}$ that $\mathsf{M}_{\text{cpl}}$ can transition to will contain $x_{0}^{\prime}$ ($\exists s_{0}^{\prime}\in S_{0}^{\prime}$ such that $x_{0}^{\prime}\in s_{0}^{\prime}$). So starting from state $x$, the set of distributions of the next state defined by $\mathsf{M}_{\text{cpl}}||\textsf{M}_{\text{per}}$ contain the distribution of the next state for $\mathsf{M}_{\text{sys}}$ if the detection probability interval for every bin $s_{i}$ of $\textsf{M}_{\text{per}}$ contains $P_{s_{i}}(h)$.
@@
[…]
$1-\sum_{b\in S}\frac{\alpha_{MC}}{|S|}\leq 1-\alpha_{MC}$